OpenStudy (anonymous):

If a stone is thrown down at 100 ft/s from a height of 1,250 feet, its height after t seconds is given by s = 1,250 − 100t − 16t^2 Estimate its instantaneous velocity at time t = 2

OpenStudy (swissgirl):

We know that the equation for instantaneous velocity is \( f'(x)=\large \frac{f(x+h)-f(x)}{(x+h)-x}\)

OpenStudy (swissgirl):

Are you familiar with this equation?

OpenStudy (anonymous):

yes

OpenStudy (swissgirl):

so \(f(t)=1250-100t-16t^2\) \(f(t+h)=1250-100(t+h)-16(t+h)^2\) so \(f'(x)= \large \frac{(1250-100(t+h)-16(t+h)^2)-(1250-100t-16t^2)}{(x+h)-x}\) \(f'(x)=\large \frac{1250-100t-100h-16t^2-32th-16h^2-1250+100t+16t^2}{h}\) \( f'(x)=\large \frac{-100h-32th-16h^2}{h}\) \(f'(x)= \large \frac{h(-100-32t-16h)}{h}\) \(f'(x)=(-100-32t-16h)\)

OpenStudy (swissgirl):

Ok Are you familiar with limits? Like what does f'(t) equal as h approaches 0?

OpenStudy (swissgirl):

Did ya follow so far?

OpenStudy (anonymous):

Yes I follow so far

OpenStudy (swissgirl):

Ok and keep in mind by accident I used X's in stead of T's so just ignore that

OpenStudy (swissgirl):